After 15 yrs, Salma will be four times as old as she is now. Determine her present age.

Answer by lenny460(1073)

After 15 yrs, Salma will be four times as old as she is now. Determine her present age.

Let Salma's age be n years.
After 15 years, Salma will be (n + 15) years
Therefore:
n + 15 = 4n
Subtract n from each side of the equation.
n + 15 - n = 4n - n
Combine like terms:
n - n = 0
4n - n = 3n
Therefore:
15 = 3n or
3n = 15
Divide each side by 3
3n/3 = 15/3
n = 5
Therefore:
Salma is 5 years old now.
In 15 years Salma will be (5 + 15) = 20 years.



Lennox Obuong
Algebra Student
Email: obuong3@aol.com